a.) An inductor carries a current of 4.4 A, and the magnetic energy stored in the inductor is 0.0022 J. What is the inductance?
  mH

b.) An inductor with L = 1.4 mH is connected to a circuit that produces a current increasing steadily from 1.8 A to 5.2 A over a time of 2.6 s. What is the voltage across the inductor?

mV
Answer #1

Given

a)


Inducctor having carrying current I = 4.4 A

energuy stored in the inductor is U = 0.0022 J

inductance of the inductor L = ?


we kmow that the magnetic energy stored in the inductor is U = 0.5*L*I^2

   L = 2*U/I^2

   L = 2*0.0022/(4.4^2) H

   L = 0.0002272727273 H

   L = 0.22727 mH


b)

Inductance L = 1.4 mH


increasing current from 1.8 A to 5.2 A , change in current dI =(5.2-1.8) = 3.4 A
the duration of time is dt = 2.6s for the change in current

voltage (induced emf) across the inductor e = ?


we know when there is a current change in the inductor then the induced emf will be developed across the inductor which is given by


   e = -L(dI/dt)


   e = 1.4*10^-3(3.4/2.6) V

   e = 0.0018307692308 V

   e = 1.83077 m V